Lori Cook has developed the following forecasting​ model:

ModifyingAbove y with carety

​=

36.036.0

​+

4.304.30​x,

where         

ModifyingAbove y with carety

​= demand for Kool Air conditioners and

                                               x​ = the outside temperature​ (degrees Fahrenheit)

​a) When the temperature outside is

7070degrees°

​F, demand forecast​ =

nothing

air conditioners ​(enter your response as an​ integer).

​b) When the temperature outside is

8080degrees°

​F, demand forecast​ =

nothing

air conditioners ​(enter your response as an​ integer).

​c) When the temperature outside is

9090degrees°

​F, demand forecast​ =

nothing

air conditioners ​(enter your response as an​ integer).
Answer #1

a) y = 36 + 4.3*X

X is the outside temperature

When X = 70°F

y = 36 + 4.3*70

= 337 Air conditioners

b) When X = 80°F

y = 36 + 4.3*80

= 380 Air conditioners

c) When X = 90°F

y = 36 + 4.3*90

= 423 Air conditioners